# Heliometric stereo: a new frontier in surface profilometry

### v1

@@ I Introduction
Accurate measurements of surface topography (often termed surface metrology) are crucial to modern research and development, with widespread applications from understanding material behavior to optimizing manufacturing[1, 2]. Significantly, surface metrology can also provide assurance that products and materials meet required standards for functionality, quality, and safety[3]. As new technologies emerge, for example, in micro and nanoscale organic and quantum devices, reliable measurement of surface topography on an ever smaller scale is crucial. Scanning helium microscopy (SHeM)[4, 5, 6, 7, 8] is a technique that uses a narrow beam of neutral helium atoms (), generated via a collimating pinhole to produce topographic micrographs. The use of a neutral atom beam has two key advantages for topographic measurements. First, it delivers an incident low energy beam without the resolution being diffraction limited: the typical beams used in SHeM offer a de Broglie energy of , corresponding to a wavelength of , [9, 10], rendering beam damage, as observed with charged beam methods[11], impossible. Second, thermal helium atoms scatter off the outermost electron density distribution, located above the ion cores of surface atoms, which results in information exclusively from the surface, without any contribution due to beam penetration into the bulk[9]. Existing scattering techniques for performing surface metrology, such as optical profilometry, laser scanning confocal microscopy, scanning electron microscopy, interferometric profilometry and structured light scanning[12, 13, 14], either use highly energetic probe particles, which can induce changes in the sample, or can have a non-negligible interaction volume with the surface and thus do not measure the surface topography exclusively. Contact techniques, such as stylus profilometers, scanning tunneling microscopes and atomic force microscopes[15, 16, 17, 18], provide a more direct surface profile measurement through contact with a physical probe. However, these are largely limited in either lateral resolution (stylus profilometers) or in aspect ratio (AFM/STM), can induce changes in the sample due to the direct contact with the probe, and in all cases convolution with the probe tip must be considered.
@@
In this paper, we describe the development of 3D imaging using SHeM (a technique termed heliometric stereo) using both single detector [25], and multiple detector instruments. We also provide detailed analysis of heliometric stereo reconstruction accuracy in general, how it is affected by contrast mechanisms unique to SHeM, and how these sources of error can be mitigated using instrumentation, algorithms and/or careful sample selection[26]. Sections II, III and IV restate the heliometric stereo method, supporting theory and surface reconstructions using a single detector, respectively, as presented by Radić et al.[25] with more detail than originally presented.
@@ II Heliometric Stereo
Analogously to photometric stereo in optics, heliometric stereo is based on the assumption that a given point on a surface will scatter the incident helium beam with a known scattering distribution, formally termed the bidirectional reflection distribution function (BRDF)[27]. This distribution defines the scattered intensity as a function of the incoming and outgoing angles of the helium beam relative to the surface. If the microscope detector is fixed at a known angle to the surface and the incident direction is known, then the recorded intensity is solely a function of surface orientation and scattering distribution. SHeM instruments have a well-defined scattering geometry[28], which gives a direct relationship between the intensity of each pixel in a micrograph and the local surface orientation of the corresponding point on the sample.
Recent experimental results have shown that the majority of unprepared surfaces studied in SHeM exhibit almost exclusively diffuse, or close to diffuse, scattering [28, 24, 20, 19, 29, 30] enabling, crucially, the first experimental implementation of heliometric stereo[25]. These materials have typically been unprepared ‘technological samples’, in contrast to conventional surface science studies where surfaces are prepared and maintained in a pristine, atomically perfect state, and where specular or Bragg scattering is often dominant[9, 10, 31, 32, 33, 34]. The observation of diffuse scattering from technological surfaces arises from the fact that the de Broglie wavelength of the helium atoms () is comparable to the interatomic spacing in solids. As such, given that most technological samples have atomically rough surfaces, diffuse scattering dominates. Consequently, the observed diffuse distribution is consistent with Knudsen’s cosine law[35], and is analogous to Lambert’s cosine law for visible light[36].
@@ III Theory
The following derivation is a summary of the heliometric stereo method previously presented by Lambrick & Palau et al.[26] Mathematically, diffuse scattering follows a cosine distribution centered on the surface normal,
@@
The orientation of a surface can be defined by two angles, plus, following the formalism of photometric stereo, we allow for a constant of proportionality, the albedo factor ( in equation (1)), hence there are 3 degrees of freedom. Therefore, a minimum of 3 data points (detector directions) are needed for each position on the surface to reconstruct the 3D topography. In practice SHeM data is often signal-to-noise limited and there may be minor deviations to the cosine law, thus at least 4 are recommended, but as demonstrated by our results in section V, not necessarily required, for a reconstruction to be robust. Indeed, the method has proved robust to noise in simulated SHeM micrographs[26].
@@ IV Reconstruction using a single-detector instrument
The single detector SHeM at Cambridge (A-SHeM), was designed to operate with an incidence angle of [5] enabling topographic imaging as well as studies of specular scattering. For the current work we modified the A-SHeM to operate in a normal incidence configuration through a new 3D printed ‘pinhole plate’ optical element[41, 42]. The pinhole plate mounts the collimating pinhole, defines the detection geometry and, in the A-SHeM, is used to mount the sample and sample manipulation stages. The old () (a) and new () (b) configurations are shown in figure 3, with the incident helium beam and rotation axis highlighted. The new pinhole plate macroscopically changes the mounted orientation of both the sample and the sample manipulation stages relative to the incident beam: in figure 3 (b) the beam is now incident at to the sample and the axis of rotation is parallel to the beam. However, while the rotation axis and the beam are parallel, they are not coaxial, thus an alignment procedure was needed to ensure that the reconstructed image area remained centered in the micrographs as the sample was rotated. The same rotation and alignment procedure as used by von Jeinsen et al.[10] was applied here. While the the new design does introduce a small () change in the path length from the collection aperture to the detector (figure 3), the detection efficiency, and therefore image acquisition, is unaffected[20].
Panel 1 adapted from a figure by T. Piesk, CC BY 4.0, published 2018, https://commons.wikimedia.org/w/index.php?curid=66347083.
@@
The reconstruction process is illustrated in figure 4 and starts by manually identifying common points across the micrographs to correlate and align the separate azimuthal images. Distinct points on the sample were selected that are accurately identifiable across micrographs, as shown in panel 2. A transformation matrix, using only translational and rotational components, was calculated to create a mapping of each micrograph onto the same axis. Mapping to a common set of axes is crucial because the pixels in each image must correspond to the same point on the sample surface so that surface orientation may be inferred. Panel 3 shows the images after alignment, with 3 tracking points highlighted (out of a total 6 used for image alignment). Where at least 3 micrographs overlapped, the intensities from the micrographs were used to calculate the surface normals according to equation (4). Normal vectors for each pixel were generated, as shown in Figure 4 panel 4. Pixels that cannot be correlated across at least 3 micrographs are assigned a surface orientation parallel to the axis because otherwise the linear system, shown in equation (2), becomes underdetermined. Orientation parallel to the axis was chosen because it is the average of all possible orientations and thus underdetermined pixels have minimal effect on the integration of surface normal vectors which gives the final surface reconstruction. Finally, the normal map is integrated using the regularized least squares method[39, 40], giving the surface reconstruction, shown in panel 5. The reconstruction captures all qualitative aspects of the crystal: we are clearly observing the top half of a truncated octahedron.
@@
The angle between the top facet of the crystal and the side facets for a truncated octahedron are (to 3 s.f.). To measure the angle between the reconstructed facets, the surface height going left to right across the crystal and top to bottom were averaged, producing two data sets each crossing the top of a trapezoid. Figure 5 shows a surface height map of the reconstruction in figure 4, with the extracted profiles horizontally and vertically across the crystal. Linear regression was applied to each edge of the two trapezoidal plots, the regression lines shown in colour in figure 5, resulting in four measurements of the facet angle; , , and , with a mean value of a overestimate compared to the ideal octahedron. The uncertainties in the linear regression were small, therefore the deviations from the expected value are not attributed to random noise.
@@ V Reconstruction using a multiple-detector instrument
The reconstruction presented in this section marks a 6-fold decrease in acquisition time compared to a single detector implementation [25]. In the current work we also use a angle of incidence instead of normally incident beam in section IV. Non-normal incidence has significant implications for the reconstruction quality of vertical surfaces which will be discussed in detail in section VI.
@@
To demonstrate multiple-detector heliometric stereo we selected a sample with known geometry for quantitative assessment of reconstruction quality. Two metal spheres with diameters were embedded in a substrate and imaged in the B-SHeM to obtain three simultaneous images that were then used to generate a surface profile using the heliometric stereo method. Figure 7 shows the acquired SHeM micrographs and the corresponding reconstructions of the spheres.
@@
Using the known spherical geometry and dimension of the two spheres we can evaluate the quality of the heliometric stereo reconstruction via the overall root-mean-square (RMS) error, giving the total deviation of the reconstruction from the true shape. The RMS shape error is defined as the overall RMS error with the height difference contribution removed to provide a measure of the curvature of the sphere. Table 2 contains the calculated values of aspect ratios, overall RMS errors and RMS shape errors for spheres A and B, alongside the signal-to-noise ratios (SNR) for each of the three detectors.
@@
The reconstruction of sphere A was qualitatively good with an overall RMS error calculated as 11.2% which exceeds the 3.4% predicted computationally by Lambrick et al.[26] for non-normal incidence scattering conditions. The difference can be attributed to two factors: (1) we use fewer detectors in the presented work than in the computational evaluation, 3 and 4, respectively. (2) The detectors used by B-SHeM have, on average, approximately a factor of 2 poorer SNR than the simulated detection where SNR=30 was set. Lambrick et al. found that the RMS error scales approximately with the SNR, which would then predict the 6. 8% RMS error with 4 images or the 7. 9% RMS error assuming the only impact of fewer images is a reduction in SNR (best case scenario), leaving 3.3% unaccounted for.
@@
By contrast, the reconstruction for sphere B is qualitatively poorer. Indeed, Table 2 shows that the overall and shape errors are larger for sphere B. The greater size of sphere B means that it exhibits greater masking and shadowing in its SHeM micrographs and consequently decreasing reconstruction accuracy [26]. These effects are explored in more detail in the next section.

### v2

@@ I Introduction
Accurate measurements of surface topography (often termed surface metrology) are crucial to modern research and development, with widespread applications from understanding material behavior to optimizing manufacturing[1, 2]. Significantly, surface metrology can also provide assurance that products and materials meet required standards for functionality, quality, and safety[3]. As new technologies emerge, for example, in micro/nanoscale organic and quantum devices, reliable measurement of delicate surface topographies on an ever smaller scale is crucial. Scanning helium microscopy (SHeM)[4, 5, 6, 7, 8] is a technique that uses a narrow beam of neutral helium atoms (), generated via a collimating pinhole to produce topographic micrographs. The use of a neutral atom beam has two key advantages for topographic measurements. First, SHeM uses a low energy beam, with a typical de Broglie energy and wavelength of and , respectively [9, 10]. Thus rendering beam damage, as observed with charged beam methods[11], impossible. Second, thermal helium atoms scatter off the outermost electron density distribution which results in information exclusively from the surface, without any contribution due to beam penetration into the bulk[9]. The atomic scale wavelength of the helium atoms gives rise to cosine-like scattering from almost any surface[12], with the exception being atomically flat and adsorbate-free crystals[10, 13], giving the technique a distinct lack of non-topographic contrast mechanisms that place it ideally for profilometry of most any surface. Existing non-contact techniques, such as optical profilometry, laser scanning confocal microscopy, scanning electron microscopy, interferometric profilometry and structured light scanning[14, 15, 16], either use highly energetic probe particles, which can induce changes in the sample, or can have a non-negligible interaction volume with the surface and thus do not measure the surface topography exclusively. These methods also have a variety of secondary contrast mechanisms that can complicate topographic reconstruction whereas SHeM almost exclusively exhibits diffuse scattering. Contact techniques, such as stylus profilometers, scanning tunneling microscopes and atomic force microscopes[17, 18, 19, 20], provide a more direct surface profile measurement through contact with a physical probe. However, these are largely limited in either lateral resolution (stylus profilometers) or in aspect ratio (AFM/STM), can induce changes in the sample due to the direct contact with the probe, and in all cases convolution with the probe tip must be considered.
@@
In this paper, we describe the development of 3D imaging using SHeM (a technique termed heliometric stereo) using both single detector [27], and multiple detector instruments[28]. We provide detailed analysis of heliometric stereo reconstruction accuracy, how it is enabled and affected by contrast mechanisms unique to SHeM, and how these sources of error can be mitigated using instrumentation, algorithms and/or careful sample selection[29]. Sections II, III and IV restate the heliometric stereo method conceptually, the supporting theory and first experimental surface reconstructions using a single detector, respectively, as presented by Radić et al.[27] and Lambrick & Palau et al.[29]. The version presented in the current work extends the original analysis and discussion than was previously presented by Radić et al.[27]. In sections II-IV we include additional discussion and analysis on the original results that lies beyond the scope of the original manuscripts’ Letter formats, providing a complete, and in-depth, understanding of heliometric stereo such that this manuscript may be an entirely self-contained reference of the method in both theory and experiment. Previously published content is indicated in text with the relevant citation. Sections V-VI present previously published results of multi-detector reconstructions, by Zhao et al.[28], alongside new data in order to provide a comprehensive error analysis on heliometric stereo, highlighting under what circumstances the method can be applied with high accuracy.
@@ II Heliometric Stereo
Analogously to photometric stereo in optics, heliometric stereo is based on the assumption that a surface will scatter the incident particles/waves with a known scattering function, formally termed the bidirectional reflection distribution function (BRDF)[30]. The BRDF defines the scattering probability as a function of the incident direction and outgoing scattered direction of the helium atoms relative to the surface orientation. In a SHeM, if the microscope detector is fixed at a known position and angle relative to the surface and the direction of the incident helium beam is known, then the remaining variable in the BRDF describing the recorded intensity is the surface orientation. The well defined scattering geometry of SHeM[12] satisfies the above description. Therefore to perform 3D surface mapping all that is needed in addition to the recorded intensity (helium micrographs) is a description of the BRDF.
Recent experimental results have shown that the majority of unprepared surfaces studied in SHeM exhibit almost exclusively diffuse, or close to diffuse, scattering [12, 26, 22, 21, 13, 31, 32, 33]. The observation of diffuse scattering provides the model for the BRDF (generally referred to simply as the scattering distribution in SHeM) required to perform surface reconstructions. Crucially, these observations enabled the first experimental implementation of heliometric stereo[27]. These materials have typically been unprepared ‘technological samples’, in contrast to conventional surface science studies where surfaces are prepared and maintained in a pristine, atomically perfect state, and where specular or Bragg scattering is often dominant[9, 10, 34, 35, 36, 37]. The observation of diffuse scattering from technological surfaces arises from the fact that the de Broglie wavelength of the helium atoms () is comparable to the interatomic spacing in solids. As such, given that most technological samples have atomically rough surfaces, diffuse scattering dominates. Consequently, the observed diffuse distribution is consistent with Knudsen’s cosine law[38], and is analogous to Lambert’s cosine law for visible light[39].
It is important to note that for SHeM there are known exceptions to the model of diffuse scattering used in heliometric stereo, the primary exception being multiple scattering, a contribution to the detected signal that arises from more than one scattering event occurring between the source and detector apertures. Discussion on the effects of multiple scattering on contrast, and therefore reconstruction accuracy, is presented theoretically in section III, and experimentally in section VI.
@@ III Theory
Here we restate the theoretical basis of heliometric stereo as first presented by Lambrick & Palau et al.[29] with text and figures largely unchanged from the original manuscript so that the current work may serve as a complete reference for the theory and experimental implementations of heliometric stereo to-date. The foundation of the method allows us to build up the two experimental implementations in sections IV and V, while the description below is self-contained for the purposes of understanding the following experimental results, full details, including supporting calculations, can be found in Lambrick & Palau[29].
Mathematically, diffuse scattering follows a cosine distribution centered on the surface normal,
@@
The orientation of a surface can be defined by two angles, plus, following the formalism of photometric stereo, we allow for a constant of proportionality, the albedo factor ( in equation (1)), hence there are 3 degrees of freedom. The albedo factor is borrowed from light scattering and has the same meaning in the current context: the albedo factor is the proportion of scattered flux that is diffusely scattered, with the rest of the flux not following equation 1, and can therefore vary across a sample due to roughness or material, or contrast features such as masking and multiple scattering. The albedo factor, along with surface normal, is calculated per-pixel when solving the system of linear equations in 5. Therefore, a minimum of 3 data points (detector directions) are needed for each position on the surface to reconstruct the 3D topography. In practice SHeM data is often signal-to-noise limited and there may be minor deviations to the cosine law, thus at least 4 are recommended, but as demonstrated by our results in section V, not necessarily required, for a reconstruction to be robust. Indeed, the method has proved robust to noise in simulated SHeM micrographs[29].
@@ IV Reconstruction using a single-detector instrument
In this section we present previous results on the experimental implementation of single detector heliometric stereo by Radić et al.[27]. The results, and supporting discussion, on the single-detector implementation of heliometric stereo presented in this section is largely unchanged from the original manuscript[27], but included in full detail such that the current manuscript can serve as a complete record of heliometric stereo theory and its experimental implementations. Figures 3, 4, 5, and the corresponding text, have been reproduced with permission from Radic̀ et al.[27], with extensions to the analysis and discussion specifically concerning the impacts of masking, shadowing and multiple scattering.
@@
The single detector SHeM at Cambridge (A-SHeM), was designed to operate with an incidence angle of [5] enabling topographic imaging as well as studies of specular scattering. To perform heliometric-stereo reconstruction it was necessary to the A-SHeM to operate in a normal incidence configuration through a new 3D printed ‘pinhole plate’ optical element[43, 44]. The pinhole plate mounts the collimating pinhole, defines the detection geometry and, in the A-SHeM, is used to mount the sample and sample manipulation stages. The old () (a) and new () (b) configurations are shown in figure 3, with the incident helium beam and rotation axis highlighted. The new pinhole plate macroscopically changes the mounted orientation of both the sample and the sample manipulation stages relative to the incident beam: in figure 3 (b) the beam is now incident at to the sample and the axis of rotation is parallel to the beam. However, while the rotation axis and the beam are parallel, they are not coaxial, thus an alignment procedure was needed to ensure that the reconstructed image area remained centered in the micrographs as the sample was rotated. The same rotation and alignment procedure as used by von Jeinsen et al.[10] was applied here. While the the new design does introduce a small () change in the path length from the collection aperture to the detector (figure 3), the detection efficiency, and therefore image acquisition, is unaffected[22].
Panel 1 adapted from a figure by T. Piesk, CC BY 4.0, 2018, https://commons.wikimedia.org/w/index.php?curid=66347083.
@@
The reconstruction process is illustrated in figure 4 and starts by manually identifying common points across the micrographs to correlate and align the separate azimuthal images. Distinct points on the sample were selected that are accurately identifiable across micrographs, as shown in panel 2. A Euclidean (or rigid) transformation matrix, one where only translation and rotation are permitted, was calculated to create a mapping of each micrograph onto the same axis. Mapping to a common set of axes is crucial because the pixels in each image must correspond to the same point on the sample surface so that surface orientation may be inferred. Panel 3 shows the images after alignment, with 3 tracking points highlighted (out of a total 6 used for image alignment). Where at least 3 micrographs overlapped, the intensities from the micrographs were used to calculate the surface normals according to equation (4). Normal vectors for each pixel were generated, as shown in Figure 4 panel 4. Pixels that cannot be correlated across at least 3 micrographs are assigned a surface orientation parallel to the axis because otherwise the linear system, shown in equation (2), becomes underdetermined. Orientation parallel to the axis was chosen because it is the average of all possible orientations and thus underdetermined pixels have minimal effect on the integration of surface normal vectors which gives the final surface reconstruction. Finally, the normal map is integrated using the regularized least squares method[41, 42], giving the surface reconstruction, shown in panel 5. The reconstruction captures all qualitative aspects of the crystal: we are clearly observing the top half of a truncated octahedron.
@@
The angle between the top facet of the crystal and the side facets for a truncated octahedron are (to 3 s.f.). To measure the angle between the reconstructed facets, the surface height going left to right across the crystal and top to bottom were averaged, producing two data sets each crossing the top of a trapezoid. Figure 5 shows a surface height map of the reconstruction in figure 4, with the extracted profiles horizontally and vertically across the crystal. Linear regression was applied to each edge of the two trapezoidal plots, the regression lines shown in color in figure 5, resulting in four measurements of the facet angle; , , and , with a mean value of a overestimate compared to the ideal octahedron. The uncertainties in the linear regression were small, therefore the deviations from the expected value are not attributed to random noise.
@@ V Reconstruction using a multiple-detector instrument
To demonstrate multiple-detector heliometric stereo we selected a sample with known geometry for quantitative assessment of reconstruction quality. Two metal spheres with diameters were embedded by different amounts in a substrate and imaged in the B-SHeM to obtain three simultaneous images that were then used to generate a surface profile using the heliometric stereo method. Figure 7 shows the acquired SHeM micrographs and the corresponding reconstructions of the spheres.
The reconstruction presented in this section marks a decrease in acquisition time per reconstructed pixel compared to a single detector implementation [27]. The decrease in acquisition time arises from both the simultaneous acquisition of images, and the complete usage of all pixels acquired, as noted in section IV 24% of the image area is unreconstructed due to the different fields of view of the five micrographs. The single (A-SHeM) and multiple detector (B-SHeM) instruments have different detector dwell and response times. If all specific instrument parameters are ignored then the relative speed up would be which accounts only for the simultaneous acquisition of images and the 24% of pixels unused in the single detector reconstruction. A speed up of roughly is therefore achieved with the multiple detector approach similar to that presented here. We note that improvements could be made in the acquisition speed for single detector approach via alternative, non-rectangular, raster patterns, which could reduce the fraction of the image area unreconstructed; these improvements may also allow for a reduction in the number of images required. In general, and for an ideal system the speed-up would be equal to the number of detectors used in the multiple detector setup — for the B-SHeM setup compared to an equivalent single detector setup with ideal image alignment — but in practice this is a lower bound on the improvement because of mechanical drift and imperfect matching of raster patterns. Details of the imaging parameters for all images used for reconstructions is given in Supplementary Materials sections SII and SIII. In the current work, we also use a angle of incidence instead of normally incident beam in section IV. Non-normal incidence has significant implications for the reconstruction quality of vertical surfaces, which will be discussed in detail in section VI.
@@
Using the known spherical geometry and dimension of the two spheres we can evaluate the quality of the heliometric stereo reconstruction via the overall root-mean-square (RMS) error, which here is defined as the squared deviation of the reconstruction from the known geometry averaged over all pixels and normalized by the known height of the sphere:
with being the reconstructed height and the height derived from the known geometry and the height of the sphere. The average is over all pixels considered to be part of the sample and not substrate, the areas used in the averaging are given in supplementary information SIV. The RMS gives a measure of the total deviation of the reconstruction from the true shape. Following from the definition used by Lambrick & Salvador-Palau et al.[29] the RMS shape error allows the reconstructed surface heights to vary: and we enforce the known maximum height of the sample, . The parameters and are found by finding the value of for the flat substrate, and the maximum height in the reconstruction, the difference between the maximum height and the flat substrate is then required to be equal to . Note that because the scaling is based on a single known parameter, and not an optimization, the shape error is not restricted to being smaller than the RMS error, although that is often the case. The metric is useful as it separates two different errors in the reconstruction, that of the shape of the surface, and the overall height of the reconstruction. Where the shape is recovered reasonably well but the overall height is under- or over-estimated the shape error will come out smaller than the RMS error. If the shape itself is reconstructed poorly then the shape error will be comparable to the RMS error. Underestimation of the height was observed with some simulated data, when multiple scattering was included, by Lambrick & Salvador-Palau et al.[29] (figure 13) and therefore the shape error allows us to disentangle if the overall observed deviations are from this underestimation or due to false reconstructed shapes. Table 2 contains the calculated values of aspect ratios, defined here as the height of the sphere above the substrate over the diameter of the sphere, overall RMS errors and RMS shape errors for spheres A and B, alongside the signal-to-noise ratios (SNR) for each of the three detectors. We limit the calculation of the RMS and shape errors to the sample itself and exclude most of the flat substrate, details are given in the SI.
@@
The reconstruction of sphere A was qualitatively good with an overall RMS error calculated as 11.2% which exceeds the 3.4% predicted computationally by Lambrick & Salvador-Palau et al.[29] for non-normal incidence scattering conditions with 4 detectors. The difference can be attributed to two factors: (1) we use fewer detectors in the presented work than in the computational evaluation, 3 and 4, respectively. (2) The detectors used by B-SHeM have, on average, approximately a factor of 2 poorer SNR than the simulated detection where SNR=30 was set. Lambrick et al. found that the RMS error scales with the SNR. The precise scaling of the scaling of SNR and reconstruction accuracy is unclear, although the simulated efforts in figure 9 of [29] seem to show that increasing the number of detectors has a similar effect to increasing the SNR of the images. Given the scaling in figure 9 of [29] the difference in signal to noise between the simulated results and the current work is likely only a partial explainer of the difference in RMS error.
@@
By contrast, the reconstruction for sphere B is qualitatively poorer. Indeed, Table 2 shows that the scaling to produce the shape error is marginally larger than the raw RMS error, the difference is very small and we consider to be insignificant. The greater size of sphere B means that it exhibits greater masking and shadowing in its SHeM micrographs and consequently decreasing reconstruction accuracy [29]. These effects are explored in more detail in the next section.